OpenStudy (geekfromthefutur):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of three halves. y - 5 = three halves(x + 1) y + 1 = three halves(x - 5) y + 5 = three halves(x - 1) y - 1 = three halves(x + 5)

OpenStudy (anonymous):

in order to determine the equation of line passing through the point (x1,y1) and having slope m we will use this form y-y1 = m*(x-x1)

OpenStudy (geekfromthefutur):

ok

OpenStudy (geekfromthefutur):

@robtobey plz help me

OpenStudy (geekfromthefutur):

hey rebtoby are you there?

OpenStudy (anonymous):

Yes.

OpenStudy (geekfromthefutur):

oh ok jw cuse you where quite lol but hey can u help me with this math question?

OpenStudy (anonymous):

The equation of the given line is: y = 3/2 x + 17/2 Subtract 1 from each side. y - 1 = (3 x)/2+15/2 Factor the RHS above. y - 1 = (3/2) (x + 5)